Sin x +Cos x= 21/2 Sin x.
Then Cos x – Sin x=?

Grade:NULL

2 Answers

noogler
8 years ago
cosx – sinx = + 21/2cosx
noogler
8 years ago
let cosx-sinx=k, sinx +cosx= 21/2sinx
square and add both eq.            u get
1+1=2sin2x+k2 …............as sin2x+cos2x=1
solve it …..........
 
APPROVE IF USEFUL!!!!!!!!!!!!!!!